OpenStudy (anonymous):

You start out with $20 and then spend money in a store where every item is $3. Y = -3x + 20. Type how to graph the equation using the y-intercept and slope. Be sure to tell where you place the first point and how to find the location of the second point.

jimthompson5910 (jim_thompson5910):

When x is 0, y is what?

OpenStudy (anonymous):

20?

jimthompson5910 (jim_thompson5910):

so you have one point (0,20)

jimthompson5910 (jim_thompson5910):

when x is 1, y is what

OpenStudy (anonymous):

not exactly sure...

jimthompson5910 (jim_thompson5910):

y = -3x + 20 y = -3(1) + 20 y = -3 + 20 y = ???

OpenStudy (anonymous):

20

OpenStudy (anonymous):

maybe

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

-3+20 is the same as 20 - 3

OpenStudy (anonymous):

inverse operation

jimthompson5910 (jim_thompson5910):

20 - 3 = ???

OpenStudy (anonymous):

17

jimthompson5910 (jim_thompson5910):

so you have the second point (1,17)

jimthompson5910 (jim_thompson5910):

Alternatively, you can start at the point (0, 20) and move 3 units down, then one unit to the right (since slope = rise/run = -3/1) to get to the next point (1, 17)

jimthompson5910 (jim_thompson5910):

either way, you have 2 points: (0,20) and (1,17) Plot them. Then draw a straight line through them.

OpenStudy (anonymous):

thanks jim

jimthompson5910 (jim_thompson5910):

yw